OpenStudy (anonymous):

Given sin A = -.45 find angle A in degrees. Round your answer to the nearest hundredth.

ganeshie8 (ganeshie8):

you will need to use ur calculator

ganeshie8 (ganeshie8):

\(\sin A = -0.45\) \(A = \sin^{-1}(-0.45)\)

OpenStudy (anonymous):

-0.46?

ganeshie8 (ganeshie8):

thats in radians

ganeshie8 (ganeshie8):

but the question wants it in degrees

OpenStudy (anonymous):

oh okay! thank you!

ganeshie8 (ganeshie8):

\(A = \sin^{-1}(-0.45) = -0.46 ~radians = -26.74~ degrees\)

ganeshie8 (ganeshie8):

-26.74 degrees is same -26.74 + 360 degrees

ganeshie8 (ganeshie8):

that gives u 333.26 degrees

ganeshie8 (ganeshie8):

see if that makes some sense...

OpenStudy (anonymous):

i got it right! thanks

ganeshie8 (ganeshie8):

good, u wlc :)